OpenStudy (anonymous):

side lengths of 84, 85 and ___ would form a right triangle.

OpenStudy (karatechopper):

Use the Pythagorean theorem. A^2+B^2=C^2

OpenStudy (karatechopper):

84^2+85^2=c^2

OpenStudy (anonymous):

ok i can square them by just saying 84*84=7056 85*85=7225 and then i add them?

OpenStudy (karatechopper):

yes :)

OpenStudy (karatechopper):

So now what does that equal

OpenStudy (anonymous):

14281...

OpenStudy (karatechopper):

ok. Now set that equal to c^2 14281=c^2 What do you think is your next step?

OpenStudy (anonymous):

im not sure maybe divide by three? I'm terrible at this

OpenStudy (karatechopper):

don't stress it. You're learning. :) When you have a variable which is squared, equal to another number, you should take the square root. When you get to higher math levels, there will be different cases when the variable is to a higher power. But right now we are just dealing with squareds. So first lets take the square root of both c^2 and 14281

OpenStudy (karatechopper):

What will the square root of 14281 equal?

OpenStudy (anonymous):

so 14281 * itself

OpenStudy (karatechopper):

no no, take the square root of 14281